# Quasi-invariant states with uniformly bounded cocycles

### v2

@@ Abstract
We investigate the notion of quasi-invariant states introduced in [2, 3] from an analytic viewpoint. We give the structures of quasi-invariant states with uniformly bounded cocycles. As a consequence, we can apply a Theorem of Kovács and Szücs to get a conditional expectation on fixed points and another of $\o$rmer to get an invariant semifinite trace under extra assumptions.
@@ 1 Introduction
Quasi-invariant measures are a much wider class than invariant, and that is why they constitute the natural environment for the theory of dynamical systems and ergodic theory. Therefore to develop a real quantum analogue of the classical theory of dynamical systems, requires developing a theory of quasi-invariant states on general operator algebras. Recently, in [2] the authors introduced the theory of quasi-invariant states under the action of a group of *-automorphisms on a von Neumann algebra (or a *-algebra). Moreover, in [3], they provide the first example of de Finetti Theorem for a quasi-exchangeable state (which is not exchangeable). The relationship between the group $G$ and modular automorphism group, invariant subalgebras, ergodicity, modular theory, and abelian subalgebras is studied in [7]. Finally, a martingale convergence theorem for strongly quasi-invariant states is given in [8].
In this note, we wish to clarify the notion $G$-quasi-invariant state under a natural boundedness assumption. The paper is structured as follows. First, we give the structures of quasi-invariant states with uniformly bounded cocycles. We further deal with a Theorem of Kovács and Szücs. Then, we extend the main result of Størmer in [19] to the strongly quasi-invariant states with uniformly bounded cocycles.
@@ 2 Quasi-invariant states
    If $\varphi$ is $G$-strongly quasi-invariant, it holds that $x_{g}$ is strictly positive. It also holds that $x_{g}$ commutes with $x_{h}$ for all $g,h\in G$. Hence the von Neumann algebra $\mathcal{C}$ generated by $\{x_{g}:g\in G\}$ is commutative. It can be also easily shown that $\mathcal{C}$ is a subalgebra of $\mathrm{Centr}(\varphi)$, the centralizer of $\varphi$, which is defined by
    $\mathrm{Centr}(\varphi):=\{x\in\mathcal{A}:\varphi(xy)=\varphi(yx)\text{ for all }y\in\mathcal{A}\}$ (2.5)
@@
If $\mathcal{A}$ is a von Neumann algebra and $\varphi$ and $\psi$ are two states on $\mathcal{A}$. We say that $\varphi$ is $\psi$–absolutely continuous if $\psi(a^{*}a)=0$ implies to $\varphi(a^{*}a)=0$ for $a\in\mathcal{A}$. We say that $\varphi$ and $\psi$ are equivalent ($\varphi\sim\psi$) if they are absolutely continuous one with respect to the other. A state $\varphi$ is $G$-invariant (resp. $G$–quasi-inavriant) if $\varphi\circ g=\varphi$ (resp. $\varphi\circ g\sim\varphi$) for all $g\in G$. Indeed, if $\varphi$ is as in Definition 2.1, then by [[18], 5.20. Lemma, page 124] one has
$\varphi\circ g(a^{*}a)=\varphi(x_{g}a^{*}a)\leq||x_{g}||\varphi(a^{*}a),\quad\varphi(a^{*}a)\leq||x_{g}^{-1}||\varphi_{g}(a^{*}a),\quad\forall g\in G,\;\forall a\in\mathcal{A},$
@@ 3 Quasi-invariant states with uniformly bounded cocycles
We will make the following assumption
@@
$\frac{1}{\lambda}\varphi(a)\leq\varphi(x_{g}a)\leq\lambda\varphi(a)$
@@
$0\leq\varphi(g(a))=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)\leq\lambda\varphi(a)$
@@
$\Box$
@@
If $\varphi$ is a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying (A), then there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
@@
$\psi(a):=\varphi(da),\quad a\in\mathcal{A},$
@@
    For all $h,g\in G$, one has $\Gamma_{gh}=\Gamma_{h}\Gamma_{g}$.
@@
    For all $g\in G$ and $a,b\in\mathcal{A}$, $\Gamma_{g}(ab)=\Gamma_{g}(a)x_{g^{-1}}^{-1}\Gamma_{g}(b)$.
@@
    $(\Gamma_{g}(a))^{*}=x_{g^{-1}}^{-1}\Gamma_{g}(a^{*})(x_{g^{-1}})^{*}.$
Let $K$ be the $\sigma$-weak closure of $\mbox{conv}_{g\in G}\{x_{g}\}$. Since the sequence $\{x_{g}\}_{g\in G}$ is uniformly bounded, $K$ is convex and compact with respect to the $\sigma$-weak topology.
@@
Now define a continuous norm $\rho$ on $\mathcal{A}$ by $\rho(x)=\varphi(xx^{*})^{1/2}$. Let $a\neq b\in K$, one has thanks to (i), (iv), (v) and Lemma 3.1
$\rho(\Gamma_{g}(a)-\Gamma_{g}(b))^{2}$ $=$ $\varphi(\Gamma_{g}(a-b)(\Gamma_{g}(a-b))^{*})$
$=$ $\varphi(\Gamma_{g}(a-b)x_{g^{-1}}^{-1}\Gamma_{g}((a-b)^{*})(x_{g^{-1}})^{*})$
$=$ $\varphi(\Gamma_{g}((a-b)(a-b)^{*})(x_{g^{-1}})^{*})$
[…]
$\geq$ $\frac{1}{\lambda}\varphi((a-b)(a-b)^{*})>0.$
Therefore $G$ is non-contracting and by the Ryll-Nardzewski fixed point theorem [16] there exists $d\in K$ such that $\Gamma_{g}(d)=d$, i.e. $d=x_{g^{-1}}g(d)$ for all $g\in G$. Notice that for all $g\in G$, $\varphi(x_{g})=1$. Then for all $x\in K$
@@
Since $\varphi$ is a faithful state, the operator $d$ is non zero. Define
@@
This proves that $\varphi$ is $G$-quasi-invariant state with cocycle given by
@@
Since $d$ is bounded invertible operator, there exists $M>0$ such that $d^{-1}(d^{-1})^{*}\leq M$. Therefore for any $g\in G$
$g^{-1}(d^{-1}(d^{-1})^{*})\leq M\;\;\mbox{and}\;\;d^{-1}g^{-1}(d^{-1}(d^{-1})^{*})(d^{-1})^{*}\leq M^{2}.$
@@
$x_{g}\leq\lambda,\quad\forall g\in G,$
which is equivalent to
@@
Notice that for all $g\in G$, $x_{g}^{-1}=g(x_{g^{-1}})$. Then from (3.2) it follows that
$x_{g^{-1}}\geq\frac{1}{\lambda},\quad\forall g\in G$
and
$x_{g}\geq\frac{1}{\lambda},\quad\forall g\in G.$
Therefore it it is clear that assumption (A) is equivalent to
@@
As a consequence of Theorem 3.2, we prove the following
@@
Remember that from the proof of Theorem 3.2, $K$ is the $\sigma$-weak closure of $\mbox{conv}_{g\in G}\{x_{g}\}$ and the cocycle $x_{g}$’s satisfy (3.3), it follows that $d$ satisfies (3.3) and $d$ is a positive invertible operator in $\mathcal{A}$. Notice that in this case $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$ which is equivalent to $[d,\;g(d)]=0$.
@@
for some normal state $\varphi$ on $\mathcal{A}$, where $d$ is a positive invertible operator in $\mathcal{A}$ such that for all $g\in G$, $[d,\;g(d)]=0$. Then it is clear that $\varphi$ is a normal faithful $G$–strongly quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$. Since $d$ is a positive invertible operator, it is clear that the cocycles $x_{g}^{\prime}$s satisfy (A). $\Box$
4 Unitary transformations associated to quasi–invariant states
When $\varphi$ is $G$-strongly quasi-invariant state on a von Neumann algebra $\mathcal{A}$ (for a given action) with cocycle $x_{g}$’s, it is proved in [2] that there exists a unique unitary representation $U$ of $G$ on $\mathcal{H}$ implementing $\pi$. We will show in this section that for $G$-quasi-invariant states, one can always spatially implement the action of $G$ (but we don’t know if the map $g\mapsto U_{g}$ is a representation).
@@ 4 Unitary transformations associated to quasi–invariant states
We will use Haagerup formalism in this section for the standard form of $\mathcal{A}$ as in section II of [22], this will avoid heavy formulas. We denote by $\mathcal{N}$ the crossed product $\mathcal{A}\rtimes\sigma^{\varphi}$ of $\mathcal{A}$ by the modular group $\sigma^{\varphi}$ associated with $\varphi$. We denote by $\theta$ the dual action of $\mathbb{R}$ on $\mathcal{N}$. Then $\theta$ satisfies $\theta_{s}(x)=x\;\;(s\in\mathbb{R})$ iff $x\in\mathcal{A}$. Let $\tau$ be the trace on $\mathcal{N}$ such that $\tau\circ\theta_{s}=e^{-s}\tau$ and denote by $L_{0}(\mathcal{N},\;\tau)$ the space of $\tau$-measurable elements $x\in\mathcal{N}$. Define for $p=1,2$
@@
There is an identification of $\mathcal{A}_{*}$ with $L_{1}(\mathcal{A})$ assigning a density $D_{\gamma}\in L_{1}(\mathcal{A})$ to a normal linear form $\gamma\in\mathcal{A}_{*}$ (Proposition 15 [22]). The duality is defined through the use of the trace $\mbox{Tr}:L_{1}(\mathcal{A})\to\mathbb{C}$ with
$\langle D_{\gamma},\;a\rangle_{L_{1}(\mathcal{A}),\;\mathcal{A}}=\hbox{\rm{Tr}}(D_{\gamma}a)=\gamma(a).$
Since $\varphi$ is faithful state on $\mathcal{A}$, its density $D_{\varphi}\in L_{1}(\mathcal{A})$ satisfies supp$(D_{\varphi})=1$ (Lemma 2, [22]). Note that by duality $G$ acts on $L_{1}(\mathcal{A})$ and we have
@@
The space $L_{2}(\mathcal{A})$ is naturally a Hilbert space with scalar product $\langle A,B\rangle=\hbox{\rm{Tr}}(A^{*}B)$. Moreover, by Theorem 36 in [22], the GNS representation of $(\mathcal{A},\;\varphi)$ is the triple $(\mathcal{A}\subset\mathcal{N},\;L_{2}(\mathcal{A}),\;D_{\varphi}^{1/2})$.
@@
$\varphi(g(a))=\hbox{\rm{Tr}}(D_{\varphi\circ g}\,a)=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)=||x_{g}||\,\hbox{\rm{Tr}}(D_{\varphi}^{1/2}aD_{\varphi}^{1/2}).$
By Proposition 4 [22], we have $0\leq D_{\varphi\circ g}\leq\|x_{g}\|D_{\varphi}$. As $D_{\varphi}$ has full support, there exists a positive element $a_{g}$ of $\mathcal{N}$ such that
@@
Considering the dual action yields that $\theta_{s}(a_{g})=a_{g}$ for all $s\in\mathbb{R}$ and $a_{g}$ is a positive operator in $\mathcal{A}$. Now, notice that for any positive operator $a$ in $\mathcal{A}$
$\varphi(a)=\varphi\circ g(x_{g}^{-1}a)\leq||x_{g}^{-1}||\,\varphi\circ g(a).$
Hence one gets
@@
$\langle U_{g}(xD_{\varphi}^{1/2}),\,U_{g}(yD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}$ $=$ $\hbox{\rm{Tr}}(a_{g}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y)D_{\varphi}^{1/2}a_{g})$
$=$ $\hbox{\rm{Tr}}(D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y))$
@@
$=$ $\langle xD_{\varphi}^{1/2},\,yD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}.$
Since $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$, $U_{g}$ can be extended to an isometry on $L_{2}(\mathcal{A})$. Now, let $h\in L_{2}(\mathcal{A})\cap\mbox{Ran}(U_{g})$. Then, for all $a\in\mathcal{A}$, $\hbox{\rm{Tr}}(a_{g}h^{*}g^{-1}D_{\varphi}^{1/2})=0$. Notice that $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$ and $a_{g}$ is invertible. Therefore, one gets $a_{g}h^{*}=0$ and $h=0$. Hence, $U_{g}$ is a unitary operator on $L_{2}(\mathcal{A})$.
@@
$\langle U_{g}^{*}xU_{g}(yD_{\varphi}^{1/2}),\,zD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}$ $=$ $\langle xg^{-1}(y)D_{\varphi}^{1/2}a_{g},\;g^{-1}(z)D_{\varphi}^{1/2}a_{g}\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}$
$=$ $\hbox{\rm{Tr}}(a_{g}D_{\varphi}^{1/2}(g^{-1}(y))^{*}x^{*}g^{-1}(z)D_{\varphi}^{1/2}a_{g})$
$=$ $\varphi(y^{*}g(x)^{*}z))$
$=$ $\langle g(x)yD_{\varphi}^{1/2},\;zD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})},$

### v3

@@ Abstract
We investigate the notion of quasi-invariant states introduced in [2] from an analytic viewpoint. We give the structures of quasi-invariant states with uniformly bounded cocycles. As a consequence, we can apply a Theorem of Kovács and Szücs to get a conditional expectation on fixed points and another of $\o$rmer to get an invariant semifinite trace under extra assumptions.
@@ 1 Introduction
Quasi-invariant measures are a much wider class than invariant, and that is why they constitute the natural environment for the theory of dynamical systems and ergodic theory. Therefore to develop a real quantum analogue of the classical theory of dynamical systems, requires developing a theory of quasi-invariant states on general operator algebras. Recently, in [2] the authors introduced the theory of quasi-invariant states under the action of a group of *-automorphisms on a von Neumann algebra (or a *-algebra). The relationship between the group $G$ and modular automorphism group, invariant subalgebras, ergodicity, modular theory, and abelian subalgebras is studied in [6]. Moreover, a martingale convergence theorem for strongly quasi-invariant states is given in [7]. Finally, in [15], quasi-invariant actions are used in the setting of discrete quantum groups. In particular, the main technical results in [15] is the computation of the Radon-Nikodym cocycle for the action of the dual of $SUq(2)$ on its Martin boundary with respect to the canonical harmonic state defined by a quantum random walk.
In this note, we wish to clarify the notion of $G$-quasi-invariant state under a natural boundedness assumption. The paper is structured as follows. First, we give the structure of quasi-invariant states with uniformly bounded cocycle. Moreover, for $G$-quasi-invariant states, we show that one can always spatially implement the action of $G$. We further deal with a Theorem of Kovács and Szücs. Then, we extend the main result of Størmer in [20] to strongly quasi-invariant states with uniformly bounded cocycles.
@@ 2 Quasi-invariant states
    If $\varphi$ is $G$-strongly quasi-invariant, it holds that $x_{g}$ is positive invertible. It also holds that $x_{g}$ commutes with $x_{h}$ for all $g,h\in G$. Hence the von Neumann algebra $\mathcal{C}$ generated by $\{x_{g}:g\in G\}$ is commutative. It can be also easily shown that $\mathcal{C}$ is a subalgebra of $\mathrm{Centr}(\varphi)$, the centralizer of $\varphi$, which is defined by
    $\mathrm{Centr}(\varphi):=\{x\in\mathcal{A}:\varphi(xy)=\varphi(yx)\text{ for all }y\in\mathcal{A}\}.$ (2.5)
Now we introduce the following lemma which plays a crucial rule later.
Lemma 2.2 ([19], 5.20. Lemma)
Let $\varphi$ be a positive form on a $C^{*}$-algebra $\mathcal{A}$ and $a\in\mathcal{A}$. If the linear form $L_{a}\varphi$ defined on $\mathcal{A}$ by $L_{a}\varphi(x)=\varphi(ax)$ is positive, then
$L_{a}\varphi(x)\leq||a||\varphi(x),$
for all positive element $x\in\mathcal{A}$.
@@
If $\mathcal{A}$ is a von Neumann algebra and $\varphi$, $\psi$ are two states on $\mathcal{A}$, we say that $\varphi$ is $\psi$–absolutely continuous if $\psi(a^{*}a)=0$ implies to $\varphi(a^{*}a)=0$ for $a\in\mathcal{A}$. We say that $\varphi$ and $\psi$ are equivalent ($\varphi\sim\psi$) if they are absolutely continuous one with respect to the other. A state $\varphi$ is $G$-invariant (resp. $G$–quasi-inavriant) if $\varphi\circ g=\varphi$ (resp. $\varphi\circ g\sim\varphi$) for all $g\in G$. Indeed, if $\varphi$ is as in Definition 2.1, then by Lemma 2.2 one has
$\varphi\circ g(a^{*}a)=\varphi(x_{g}a^{*}a)\leq||x_{g}||\varphi(a^{*}a),\quad\varphi(a^{*}a)\leq||x_{g}^{-1}||\varphi\circ g(a^{*}a),\quad\forall g\in G,\;\forall a\in\mathcal{A},$
@@ 3 Quasi-invariant states with uniformly bounded cocycles
We will make the following assumption.
@@
$\frac{1}{\lambda}\varphi(a)\leq\varphi(x_{g}a)\leq\lambda\varphi(a)\quad\textrm{and}\quad\frac{1}{\lambda}\varphi(a)\leq\varphi(a(x_{g}^{-1})^{*})\leq\lambda\varphi(a).$
@@
$0\leq\varphi(g(a))=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)\leq\lambda\varphi(a).$
@@
Since $\varphi_{g}=\varphi\circ g:a\mapsto\varphi(x_{g}a)$ is a state, the linear form $a\mapsto\varphi(a(x_{g}^{-1})^{*})=\varphi_{g}(x_{g}^{-1}a(x_{g}^{-1})^{*})$ is positive. Then, for all $a\in\mathcal{A}$, we have
$\varphi(a(x_{g}^{-1})^{*})=\overline{\varphi(a^{*}(x_{g}^{-1})^{*})}=\overline{\varphi((x_{g}^{-1}a)^{*})}=\varphi(x_{g}^{-1}a)=L_{x_{g}^{-1}}\varphi(a)$
thanks to the positivity of the above-mentioned linear form and $\varphi$, and we can use Lemma 2.2 again as $L_{x_{g}^{-1}}\varphi\geq 0$. Similarly $\varphi=L_{x_{g}}(L_{x_{g}^{-1}}\varphi)\geq 0$ gives the last inequality. $\Box$
@@
If $\varphi$ is a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying (A), then there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded operator $d\in\mathcal{A}$ such that
$\psi(a)=\varphi(da),\quad a\in\mathcal{A}.$
@@
$\psi(a)=\varphi(da),\quad a\in\mathcal{A},$
@@
    For all $h,g\in G$, one has $\Gamma_{gh}=\Gamma_{g}\Gamma_{h}$.
@@
    For all $g\in G$ and $a,b\in\mathcal{A}$, $\Gamma_{g}(ab)=\Gamma_{g}(a)(x_{g^{-1}})^{-1}\Gamma_{g}(b)$.
@@
    $(\Gamma_{g}(a))^{*}=(x_{g^{-1}})^{-1}\Gamma_{g}(a^{*})(x_{g^{-1}})^{*}.$
Let $K$ be the $\sigma$-weak closure of $\mbox{conv}\{x_{g}\}_{g\in G}$. Since the set $\{x_{g}\}_{g\in G}$ is uniformly bounded, $K$ is convex and compact with respect to the $\sigma$-weak topology.
@@
Now define a continuous norm $\rho$ on $\mathcal{A}$ by $\rho(x)=\varphi(xx^{*})^{1/2}$. Let $a\neq b\in K$, thanks to the second estimate in Lemma 3.1 applied to $\Gamma_{g}(a-b)(\Gamma_{g}(a-b))^{*}\geq 0$ for $g^{-1}$:
$\rho(\Gamma_{g}(a)-\Gamma_{g}(b))^{2}\geq\frac{1}{\lambda}\varphi(\Gamma_{g}(a-b)(\Gamma_{g}(a-b))^{*}((x_{g^{-1}})^{-1})^{*}).$
Using the algebraic identities (v), (iv) and (iii)
$\lambda\rho(\Gamma_{g}(a)-\Gamma_{g}(b))^{2}$$\geq$ $\varphi(\Gamma_{g}(a-b)(x_{g^{-1}})^{-1}\Gamma_{g}((a-b)^{*}))$
$=$ $\varphi(\Gamma_{g}((a-b)(a-b)^{*}))$
$=$ $\varphi((a-b)(a-b)^{*})>0.$
Therefore $\Gamma_{G}$ is non-contracting and by the Ryll-Nardzewski fixed point theorem [17] there exists $d\in K$ such that $\Gamma_{g}(d)=d$, i.e. $d=x_{g^{-1}}g(d)$ for all $g\in G$. Notice that for all $g\in G$, $\varphi(x_{g})=1$. Then for all $x\in K$
@@
It follows that the operator $d$ is non zero. Define
@@
This proves that $\varphi$ is $G$-quasi-invariant state with cocycle
@@
Since $d$ is a bounded invertible operator, then for any $g\in G$
$g^{-1}(d^{-1}(d^{-1})^{*})\leq||d^{-1}||^{2}\;\;\mbox{and}\;\;dg^{-1}(d^{-1}(d^{-1})^{*})d^{*}\leq||d||^{2}||d^{-1}||^{2}.$
@@
$x_{g}\leq\lambda,\quad\forall g\in G.$
Thanks to (2.3), this is equivalent to
@@
Therefore it is clear that assumption (A) is equivalent to
@@
As a consequence of Theorem 3.2, we prove the following.
@@
Remember that from the proof of Theorem 3.2, $K$ is the $\sigma$-weak closure of $\mbox{conv}\{x_{g}\}_{g\in G}$ and the cocycle $x_{g}$’s satisfy (3.3), it follows that $d$ satisfies (3.3) and $d$ is a positive invertible operator in $\mathcal{A}$. Notice that in this case $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$ which is equivalent to $[d,\;g(d)]=0$.
@@
for some normal state $\varphi$ on $\mathcal{A}$, where $d$ is a positive invertible operator in $\mathcal{A}$ such that for all $g\in G$, $[d,\;g(d)]=0$. Then it is clear that $\varphi$ is a normal faithful $G$–strongly quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$. Since $d$ is a positive invertible operator, it is clear that the cocycle $x_{g}^{\prime}$s satisfies (A). $\Box$
4 Unitary implementation associated to quasi–invariant states
When $\varphi$ is $G$-strongly quasi-invariant state on a von Neumann algebra $\mathcal{A}$ (for a given action) with cocycle $x_{g}$’s, it is proved in [2] that, in the GNS representation of $\varphi$, there exists a unique unitary representation of $G$ implementing the action by conjugation. We will recover it in this section and show that for $G$-quasi-invariant states, one can always spatially implement the action of $G$ (but we don’t know if the map $g\mapsto U_{g}^{*}$ is a representation). Note that in [10], it is showed that the group $G$ has always a canonical unitary implementation which is different to the one introduced here. In particular in [Theorem 3.2, [10]], with the usual notation from there, the condition $u_{g}J=Ju_{g}$ is not satisfied in our case.
@@ 4 Unitary implementation associated to quasi–invariant states
We will use Haagerup formalism in this section for the standard form of $\mathcal{A}$ as in section II of [23], this will avoid heavy formulas. We denote by $\mathcal{N}$ the crossed product $\mathcal{A}\rtimes_{\sigma^{\varphi}}\mathbb{R}$ of $\mathcal{A}$ by the modular group $\sigma^{\varphi}$ associated with $\varphi$. We denote by $\theta$ the dual action of $\mathbb{R}$ on $\mathcal{N}$. Then $\theta$ satisfies $\theta_{s}(x)=x\;\;(s\in\mathbb{R})$ iff $x\in\mathcal{A}$. Let $\tau$ be the trace on $\mathcal{N}$ such that $\tau\circ\theta_{s}=e^{-s}\tau$ and denote by $L_{0}(\mathcal{N},\;\tau)$ the space of $\tau$-measurable operators $x$ affiliated with $\mathcal{N}$. Define for $p=1,2$
@@
There is an identification of $\mathcal{A}_{*}$ with $L_{1}(\mathcal{A})$ assigning a positive operator $D_{\gamma}\in L_{1}(\mathcal{A})$ to a normal linear form $\gamma\in\mathcal{A}_{*}$ (Proposition 15 [23]). The duality is defined through the use of a trace-like functional $\mbox{tr}:L_{1}(\mathcal{A})\to\mathbb{C}$ with
$\langle D_{\gamma},\;a\rangle_{L_{1}(\mathcal{A}),\;\mathcal{A}}=\mbox{tr}(D_{\gamma}a)=\gamma(a).$
Since $\varphi$ is faithful state on $\mathcal{A}$, its density $D_{\varphi}\in L_{1}(\mathcal{A})$ satisfies supp$(D_{\varphi})=1$ by Proposition 4(2c) in [23]. Note that, by duality, $G$ acts on $L_{1}(\mathcal{A})$ and we have
@@
The space $L_{2}(\mathcal{A})$ is naturally a Hilbert space with scalar product $\langle A,B\rangle=\mbox{tr}{(A^{*}B)}$. Moreover, by Theorem 36 in [23], the GNS representation of $(\mathcal{A},\;\varphi)$ is the triple $(\mathcal{A}\subset\mathcal{N},\;L_{2}(\mathcal{A}),\;D_{\varphi}^{1/2})$.
@@
$\varphi(g(a))=\mbox{tr}{(D_{\varphi\circ g}\,a)}=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)=||x_{g}||\,\mbox{tr}{(D_{\varphi}^{1/2}aD_{\varphi}^{1/2})}.$
By Theorem 7 [23], we have $0\leq D_{\varphi\circ g}\leq\|x_{g}\|D_{\varphi}$. As $D_{\varphi}$ has full support, there exists a positive element $a_{g}$ of $\mathcal{N}$ such that
@@
Considering the dual action yields that $\theta_{s}(a_{g})=a_{g}$ for all $s\in\mathbb{R}$ and $a_{g}$ is a positive operator in $\mathcal{A}$. Now, notice that by the proof of Lemma 3.1, $\varphi\leq||x_{g^{-1}}||\,\varphi\circ g$. Hence one gets
@@
$\langle U_{g}(xD_{\varphi}^{1/2}),\,U_{g}(yD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A})}$ $=$ $\mbox{tr}{(a_{g}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y)D_{\varphi}^{1/2}a_{g})}$
$=$ $\mbox{tr}{(D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y))}$
@@
$=$ $\langle xD_{\varphi}^{1/2},\,yD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}.$
Since $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$, $U_{g}$ can be extended to an isometry on $L_{2}(\mathcal{A})$. Now, let $h\in L_{2}(\mathcal{A})\cap\mbox{Ran}(U_{g})^{\bot}$. Then, for all $a\in\mathcal{A}$
$0=\langle h,\,U_{g}(aD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A})}=\mbox{tr}{(h^{*}g^{-1}(a)D_{\varphi}^{1/2}a_{g})}=\mbox{tr}{(a_{g}h^{*}g^{-1}(a)D_{\varphi}^{1/2})}.$
Notice that $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$ and $a_{g}$ is invertible. Therefore, one gets $a_{g}h^{*}=0$ and $h=0$. Hence, $U_{g}$ is a unitary operator on $L_{2}(\mathcal{A})$.
@@
$\langle U_{g}^{*}xU_{g}(yD_{\varphi}^{1/2}),\,zD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}$ $=$ $\langle xg^{-1}(y)D_{\varphi}^{1/2}a_{g},\;g^{-1}(z)D_{\varphi}^{1/2}a_{g}\rangle_{L_{2}(\mathcal{A})}$
$=$ $\mbox{tr}{(a_{g}D_{\varphi}^{1/2}(g^{-1}(y))^{*}x^{*}g^{-1}(z)D_{\varphi}^{1/2}a_{g})}$
$=$ $\varphi(y^{*}g(x)^{*}z)$
$=$ $\langle g(x)yD_{\varphi}^{1/2},\;zD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})},$